OpenStudy (anonymous):

ben has 4 more than twice as many cd's as jake. if they have a total of 31 cd's, home cd's does jake have ?

OpenStudy (anonymous):

Let Jake's number of CDs be \(x\). Then Ben has \(2x+4\) CDs. Then, what's the sum in terms of \(x\)?

OpenStudy (anonymous):

Jake = j Ben = b ben has 4 more than twice jake b = 4+2j total 31 cds b + j =31

OpenStudy (anonymous):

you can use subsitution to solve b = 4+2j b + j =31 subsitute b = 4+2j into b + j =31 (4+2j)+j=31 simplify to get 4+3j=31 3j=27 j=9 Jake has 9 CDs b=4+2j or b=4+2(9) so Ben has 22 CDs

OpenStudy (anonymous):

thankk youu! :)

OpenStudy (anonymous):

welcome, glad to help